2.4.80-T : Question Help Calculate the range, variance, and standard deviation for the following samples. a, 6, 2, 3, 0, 3 b. 6,1,4,4,7,0,7 C.1,-1,2. 4, 3, 5, 5, 2,4 d. 0, 4, 0, 0, -2,2, -4, 2, 0, -2,2, -2, 0, -5, -4, -2,0,2 a. The range is Type an integer or a decimal. Do not round.)
Answer #1

a) From the given data

S.NO. X (X-Xbar)^2
1 6 10.24
2 2 0.64
3 3 0.04
4 0 7.84
5 3 0.04
Total: 14 18.8

Range = maximum-minimum-6-0 = 6 1 14 =-=2.8 = Σ(n--)=-18.8=4.7 - SD = /variance = J47 = 2.1679

b) From the given data

S.NO. X (X-Xbar)^2
1 6 3.4488
2 1 9.8778
3 4 0.0204
4 4 0.0204
5 7 8.163
6 0 17.1636
7 7 8.163
Total: 29 46.857

Range-maximum -minimum -7-0-7 =_ = 4, 1429 - Variance(s2)--Ty( -r) = 646.857 = 7.809524 SD-/variance-ง่า 809524-27945

c) From the given data

S.NO. X (X-Xbar)^2
1 1 3.1606
2 -1 14.2718
3 2 0.605
4 4 1.4938
5 3 0.0494
6 5 4.9382
7 5 4.9382
8 2 0.605
9 4 1.4938
Total: 25 31.5558

Range maximum -minimum-5-(-1) 6 1*2:5 Sample Variance(,)-קל(h-7-831 5558-3 9445 SD- vvari ance .94445-1.9861

d) From the given data

S.NO. X (X-Xbar)^2
1 0 0.25
2 4 20.25
3 0 0.25
4 0 0.25
5 -2 2.25
6 2 6.25
7 -4 12.25
8 2 6.25
9 0 0.25
10 -2 2.25
11 2 6.25
12 -2 2.25
13 0 0.25
14 -5 20.25
15 -4 12.25
16 -2 2.25
17 0 0.25
18 2 6.25
Total: -9 100.5

4-(-5) = 9 Range = maximum-minimum i-1 Saple Vanance(,)-.-15(hx) = 171005-59118 SD = /variance = /591 18 = 2.4314 -Σ(x-x), 17100.5-5.9118